OpenStudy (anonymous):

A 12 foot board is cut into two pieces. Twice the length of shorter piece is 3ft less than the length of the longer piece. Find the lenght of each piece.

OpenStudy (anonymous):

let l = x & b = y x +y = 12 (i) as per equation ; 2x = y-3 => y = 2x +3 Putting the value of y in (i) 3x = 9 => x = 3 y = 2x 3 +3 = 9

OpenStudy (anonymous):

Thanks a lot